# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ III Congestion Mitigation Methodology
Figure 4: With respect to the UAS $i\in\{1,2,3,4\}$, the set $\{\ j^{\prime}\ |\ j\in\{1,2,3,4\}\backslash\{i\}\ \}$ are the $L$ slot look-ahead predicted positions of other UAS. When $M=2$, from the perspective of UAS 2 present in zone $Z$, zone $Z^{\prime}$ will not be congested in the next $L$ time slots as only one predicted position (that of UAS 1) exists in $Z^{\prime}$. From the perspective of UAS 3 present in zone $Z$, the zone $Z^{\prime}$ will be congested as two predicted positions (UAS 1 and 2) lie in zone $Z^{\prime}$.
The congestion definition introduced for the nominal zone applies to all the zones in the grid. However, the decisions of UAS present in an arbitrary zone $Z=(X,Y)$ are only influenced by congestion occurrences in one or two of its neighboring upstream zones, which are,
@@
    $\{(X,Y+1)\}$ if $X=0$
@@
    $\{(X-1,Y+1),(X,Y+1)\}$ if $X>0$
@@
    $\{(X,Y+1),(X+1,Y+1)\}$ if $X<0$
For ease of notation, we refer to the neighboring zones of interest as shown in Fig. 5. $Z_{\mathrm{UP}}$ is an upstream neighbor, $Z_{\mathrm{IN}}$ is an inner neighbor (zone closer to the Stream 00 zone than zone $Z$), $Z_{\mathrm{UD}}$ is an upstream inner diagonal neighbor and, $Z_{\mathrm{ON}}$ is an outer neighbor (zone farther from Stream 00 than zone $Z$). Note that $Z_{\mathrm{IN}}$ and $Z_{\mathrm{UD}}$ are defined only when $X\neq 0$. The zone notation can be extended to respective nodes as well. By interconnecting the $\widetilde{Z}=(X,Y)$ nodes with respective $\widetilde{Z}_{\mathrm{UP}}$ nodes, $X\in\{-X_{e},...,0,...X_{e}\}$, we have $12X_{e}+$ 1 parallel segments referred to as Stream $X$ segments.
Figure 5: Zone neighborhood notation for $Z=(X,Y)$ zone. The congestion occurrences in upstream zones $Z_{\mathrm{UP}}$ and $Z_{\mathrm{UD}}$ influence decisions of UAS present in $Z$.
At node $\widetilde{Z}$, if the UAS decides to avoid the zone $Z^{\prime}$, it generates a piece-wise re-route path (a sequence of the number of transitions to take in a specific heading direction). The UTM provides a heading-reference directed graph to the UAS present in an arbitrary grid level as shown in Fig. 6. For a zone $Z=(X,Y),\ X\neq 0$, the graph comprises of
@@
    an $\alpha_{Z}$ path of $S$ cells in the upstream direction (hereafter referred as $\alpha_{Z}$ heading) connecting node $\widetilde{Z}$ with $\widetilde{Z}_{\mathrm{UP}}$,
@@
    a $\widehat{\delta}_{Z}$ path of $S$ cells in the inward diagonal direction ($\widehat{\delta}_{Z}$ heading) connecting node $\widetilde{Z}$ with $\widetilde{Z}_{\mathrm{UD}}$,
  3. 3.
    a $\gamma_{Z}$ and a $\beta_{Z}$ path of $L$ cells each in the outward lateral direction ($\gamma_{Z}\rightarrow\beta_{Z}$ heading), connecting zone $Z_{\mathrm{IN}}$ with node $\widetilde{Z}$ and node $\widetilde{Z}$ with zone $Z_{\mathrm{ON}}$, respectively. These are alternate paths for UAS in $Z$, in case the $Z_{\mathrm{UP}}$ turns congested. The paths including node $\widetilde{Z}$ are collectively referred to as $\gamma_{Z}\rightarrow\beta_{Z}$ path.
  4. 4.
    $L$ number of $\widehat{\gamma}_{Z}$ inward diagonal ($\widehat{\gamma}_{Z}$ heading) and $L$ number of $\widehat{\beta}_{Z}$ outward diagonal ($\widehat{\beta}_{Z}$ heading) parallel paths, connecting ${\gamma}_{Z}$ and ${\beta}_{Z}$ path with $\alpha_{Z}$ path, respectively. These are alternate paths for UAS in the $2L$ cells of $\gamma_{Z}$ and $\beta_{Z}$ path, in case the $Z_{\mathrm{UP}}$ turns uncongested.
Figure 6: Heading reference directed graph for UAS present in an arbitrary level.
As a special case, the graph for Stream 00 zones comprises an $\alpha_{0}$ upstream path, two $\beta_{0}$ paths with opposite headings directed laterally outward from $(0,Y)$ node, and $2L$ number of $\widehat{\beta}_{0}$ inward diagonal paths. For notation convenience, we use $\alpha_{.},\beta_{.},\gamma_{.},\widehat{\beta}_{.},\widehat{\gamma}_{.},\widehat{\delta}_{.}$, with subscript .$=\mathrm{I},\mathrm{P},\mathrm{O},\mathrm{D}$ referring to the paths in the zones $Z_{\mathrm{IN}},Z_{\mathrm{UP}},Z_{\mathrm{ON}},Z_{\mathrm{UD}}$ respectively. We ignore the subscript when referring to an arbitrary zone.
The rule-based congestion mitigation strategy for the UAS foreseeing congestion in its $Z_{\mathrm{UP}}$ and $Z_{\mathrm{UD}}$ zones is as follows. The following rules assume the UAS is transitioning on a Stream $X$ segment, $X=\{0,...,X_{e}\}$. The heading reference graph is symmetric about the Stream 00 segment; hence similar rules when $X=\{-X_{e},...,-1\}$ do hold.
Rule 1.
The UAS present in zone $Z=(X,Y)$ identify respective $Z_{\mathrm{UP}}$ and $Z_{\mathrm{UD}}$ zones as congested or uncongested following Definition 2.
Note that the UAS is restricted to transition on the heading-constrained graph shown in Fig. 6. Under the above structural constraint, among the UAS present in the zone $Z$, the $L$ slot look-ahead predicted positions of only those UAS transitioning on $\alpha_{Z},\widehat{\beta}_{Z},\widehat{\gamma}_{Z}$ paths lie in zone $Z_{\mathrm{UP}}$ and are responsible for congestion. Similarly, predicted positions of UAS transitioning on $\widehat{\delta}_{Z}$ path lie in zone $Z_{\mathrm{UD}}$ (refer Fig. 7).
Figure 7: The $L=5$ slot look-ahead predicted positions of UAS, which are restricted on the heading reference graph. If $M=2$, then by Definition 2, in the perspective of UAS present at node $\widetilde{Z}$, both zones $Z_{\mathrm{UP}}$ and $Z_{\mathrm{UD}}$ are congested (due to UAS 3, 4, 5, and 7). UAS present on $\gamma\rightarrow\beta$ path (UAS 1 and 2) do not contribute to congestion in $Z_{\mathrm{UP}}$ and $Z_{\mathrm{UD}}$.
Rule 2.
In a given time slot, if the UAS present in $\gamma_{Z}\rightarrow\beta_{Z}$ path finds zone $Z_{\mathrm{UP}}$ congested, then the UAS executes one lateral transition in the $\gamma_{Z}\rightarrow\beta_{Z}$ heading direction.
In Fig. 8, if $Z_{\mathrm{UP}}$ is congested in timeslot $k$, then in timeslot $k+1$ UAS in positions $1,6,9$ would be in positions $2,7,10$ respectively. A zone congested at one timeslot may or may not be congested in the next. Only myopic position predictions ($1$ slot look-ahead) are possible for UAS transitioning on $\gamma_{Z}\rightarrow\beta_{Z}$, as it is uncertain whether these UAS would continue transitioning in $\gamma_{Z}\rightarrow\beta_{Z}$ direction or not. Hence, these UAS are excluded in the Definition 2. If the UAS that has entered the $\beta_{Z}$ path (UAS in position $7$ in Fig. 8) finds the $Z_{\mathrm{UP}}$ zone congested for $L$ successive slots, the UAS ends in the $\gamma_{\mathrm{O}}$ path. Similarly, if a UAS that has entered the $\gamma_{Z}$ path (UAS in position $1$ in Fig. 8) finds the $Z_{\mathrm{UP}}$ zone congested for $L$ successive slots, it reaches node $\widetilde{Z}$.
Figure 8: Relative position $j$ of the UAS in the $\gamma_{Z}\rightarrow\beta_{Z}$ path with respect to a fixed reference frame in the zone $Z$ ($L=5$ slots, edge length $S=2L+1=11$ timeslots).
The duration for which zone $Z_{\mathrm{UP}}$ remains congested and the number of UAS present in the zone when congested depends on the exogenous UAS in $Z_{\mathrm{UP}}$ as well. For the UAS laterally transitioning on $\gamma_{Z}\rightarrow\beta_{Z}$ path, with each successive outward transition, the UAS belief that the respective $Z_{\mathrm{UP}}$ would always remain congested is reinforced. Hence, the UAS that has transitioned longer on the $\gamma_{Z}\rightarrow\beta_{Z}$ path rather prefers to avoid $Z_{\mathrm{UP}}$ and progress to $Z_{\mathrm{ON}}$, believing that when the UAS enters $\gamma_{\mathrm{O}}$ path the upstream neighbor of $Z_{\mathrm{ON}}$ would be uncongested. The UAS that has just entered node $\widetilde{Z}$ from the downstream zones passively prefers to move towards $Z_{\mathrm{UP}}$ if uncongested. Whereas, the UAS presently in $\gamma_{Z}$ path had previously avoided the $Z_{\mathrm{UD}}$ zone believing that when it entered $\gamma_{Z}$ path the $Z_{\mathrm{UP}}$ would have been uncongested. The UAS in the $\gamma_{Z}$ path actively prefers to move towards zone $Z_{\mathrm{UP}}$, thus has a higher preference than the UAS in the node $\widetilde{Z}$ and $\beta_{Z}$ path. Each UAS present in $\gamma_{Z}\rightarrow\beta_{Z}$ path is marked with a value $\psi_{Z}\in[-L,L]$ reflecting the above time varying UAS preference to move towards $Z_{\mathrm{UP}}$ zone. The UAS preference $\psi_{Z}$ is modeled as a linear function of the UAS relative position $j$ on the $\gamma_{Z}\rightarrow\beta_{Z}$ path, given by
$\psi_{Z}=L+1-j,\ \ \ j\in\mathbb{Z}_{[1,S]}$ (2)
UAS having zero $\psi_{Z}$ would mean the UAS is currently at node $\widetilde{Z}$. UAS with negative $\psi_{Z}$ prefer reverting (inward diagonal transition) to the $Z_{\mathrm{UP}}$ zone. The UAS having positive $\psi_{Z}$ prefer progressing (outward diagonal transition) to the $Z_{\mathrm{UP}}$ zone if the zone turns uncongested. Preference $\psi_{Z}$ is information shared among UAS transitioning in $\gamma_{Z}\rightarrow\beta_{Z}$ path or can be computed using positional information $j$, enabling the UAS to distributedly decide which UAS should proceed upstream when the respective $Z_{\mathrm{UP}}$ zone turns uncongested.
Rule 3.
If in a given time slot the $Z_{\mathrm{UP}}$ zone turns uncongested, then among the UAS laterally transitioning on $\gamma_{Z}\rightarrow\beta_{Z}$ path, that UAS which has the highest preference $\psi_{Zmax}\in\mathbb{Z}_{[-L,L]}$ would proceed towards the $Z_{\mathrm{UP}}$ zone. The remaining UAS execute one lateral outward transition on respective paths.
Applying Rule 3 in Fig. 8, if UAS are present in position $1,6,9$, then UAS in position $1$ proceeds to position ‘a’, whereas UAS in position $6,9$ transition to position $7,10,$ respectively. The UAS in position $6,7$ proceed to position ‘b’ if and only if no UAS are present in positions $1-5$ and $1-6$ respectively. Corresponding to each value of $\psi_{Z}\in\mathbb{Z}_{[-L,L]}$, we have a piece-wise $\widehat{\gamma}_{Z}\rightarrow\alpha_{Z}$ or $\widehat{\beta}_{Z}\rightarrow\alpha_{Z}$ or $\alpha_{Z}$ path that connects the $\gamma_{Z}\rightarrow\beta_{Z}$ path with node $\widetilde{Z}_{\mathrm{UP}}$. Thus, a total of $2L+1=S$ paths exists. Further, the UAS in $\gamma_{Z}\rightarrow\beta_{Z}$ path needs $S$ transitions on respective $\psi_{Z}$ path to reach the $\widetilde{Z}_{\mathrm{UP}}$ node (refer Fig. 9).
Figure 9: $S$ paths connecting $\gamma_{Z}\rightarrow\beta_{Z}$ path with $\widetilde{Z}_{\mathrm{UP}}$ node.
Rule 4.
Following Rule 3, the UAS in the $\gamma_{Z}\rightarrow\beta_{Z}$ path that has the highest preference $\psi_{Zmax}$ opts the following path to relatively transition upstream and reach $\widetilde{Z}_{\mathrm{UP}}$ node.
  1. 1.
    If $\psi_{Zmax}=0$, then UAS is at node $Z$ and proceeds with $S$ successive transitions in $\alpha_{Z}$ heading direction.
  2. 2.
    When $\psi_{Zmax}<0$ then UAS proceed with $|\psi_{Zmax}|$ successive diagonal transitions in $\widehat{\beta}_{Z}$ heading direction followed by $S-|\psi_{Zmax}|$ upstream transitions in $\alpha_{Z}$ heading direction.
  3. 3.
    When $\psi_{Zmax}>0$ then UAS proceed with $|\psi_{Zmax}|$ successive diagonal transitions in $\widehat{\gamma}_{Z}$ heading direction followed by $S-|\psi_{Zmax}|$ upstream transitions in $\alpha_{Z}$ heading direction.
The following rules are for handling contingencies and exceptional scenarios. The UAS heading in $\gamma$ and ($\alpha$ or $\delta$) direction may simultaneously reach the $\widetilde{Z}$ node in some time slot $(k-1)$ and encounter conflict. If conflict is not resolved, then following rules 2-4, the UAS will thereafter always conflict with each other for cell occupancy.
Rule 5.
In a given time slot, if a UAS in $\gamma_{Z}$ path is in conflict at node $\widetilde{Z}$ with a UAS from downstream, then the UAS in $\gamma_{Z}$ path is assigned preference $\psi_{Z}>L$. Irrespective of $Z_{\mathrm{UP}}$ zone being congested, the UAS with $\psi_{Z}>L$ executes a conflict resolution maneuver (shown in Fig. 10(a)) and then proceeds with $S-1$ upstream transitions in $\alpha_{Z}$ heading direction to reach $\widetilde{Z}_{\mathrm{UP}}$.
@@
The Conflict Detection and Resolution (CDR) methods suggested in [22, 23, 24] can be applied for conflict resolution maneuvers. The UAS is assumed to be enclosed in a virtual cylindrical disk of a predefined safety radius. The cells are large enough that during the spatiotemporal resolution maneuver, the safety disks do not overlap. However, the emphasis is on the UAS at $\gamma_{Z}$ cell center encountering conflict in time instant $(k-1)\Delta T$ should reach neighboring $\alpha_{Z}$ cell center by time instant $(k+1)\Delta T$. The UAS is free to accelerate or decelerate about the nominal velocity to achieve the above time synchronization. During the conflict, the UAS executing conflict resolution maneuver (the UAS with $\psi_{Z}>L$) has the highest preference. Hence, irrespective of zone $Z_{\mathrm{UP}}$ being uncongested, the remaining UAS in $\gamma_{Z}\rightarrow\beta_{Z}$ path proceeds with one outward lateral transition.
Rule 6.
(Descend condition) In a given timeslot, if the downstream UAS enters node $\widetilde{Z}$ and finds the $Z_{\mathrm{UD}}$ zone uncongested with no UAS present in $\gamma_{\mathrm{I}}\rightarrow\beta_{\mathrm{I}}$ path, then (irrespective of $Z_{\mathrm{UP}}$ being uncongested or conflict at node $\widetilde{Z}$) the downstream UAS proceeds with $S$ successive transitions in the $\widehat{\delta}_{Z}$ heading direction (shown in Fig. 10(b)) and reaches the $\widetilde{Z}_{\mathrm{UD}}$ node.

### v2

@@ 3 Congestion Mitigation Methodology
Figure 3: In the figure, with respect to the UAS $i\in\{1,2,3,4\}$, the set $\{\ j^{\prime}\ |\ j\in\{1,2,3,4\}\backslash\{i\}\ \}$ are the $L$ slot look-ahead predicted positions of other UAS. When $M=2$, from the perspective of UAS 2 present in zone $Z$, zone $Z^{\prime}$ will not be congested in the next $L$ timeslots as only one predicted position (that of UAS 1) exists in $Z^{\prime}$. From the perspective of UAS 3 present in zone $Z$, the zone $Z^{\prime}$ will be congested as two predicted positions (UAS 1 and 2) lie in zone $Z^{\prime}$.
The congestion definition applies to all the zones in the grid. However, the decisions of UAS present in any zone $Z$ are only influenced by congestion occurrences in two of its neighboring upstream zones. These are shown as blue-colored zones in Fig. 4. In this figure, the neighborhood zones of $Z=(X,Y)$ that are of significance for the congestion mitigation strategy are denoted as $Z_{\mathrm{UP}},Z_{\mathrm{IN}},Z_{\mathrm{ID}}$ and $Z_{\mathrm{ON}}$ (notation described in Table. 1). Note that $Z_{\mathrm{IN}}$ and $Z_{\mathrm{ID}}$ are defined only when $X\neq 0$. The zone notation can be extended to respective nodes as well. At node $\widetilde{Z}$, if the UAS decides to avoid zone $Z_{\mathrm{UP}}$, it generates a piecewise re-route path (a sequence of the number of transitions in a specific heading direction).
Figure 4: Zone neighborhood of three arbitrary zones $Z=(X,Y)$ in the grid, whose $X<0$, $X=0$, and $X>0$, respectively. The congestion occurrences in respective upstream zones $Z_{\mathrm{UP}}$ and $Z_{\mathrm{ID}}$ influence the decisions of UAS present in $Z$.
Table 1: Zone neighborhood and path notation for zone $Z$
$Z_{\mathrm{UP}}$ the neighbor in the upstream direction.
$Z_{\mathrm{IN}}$ the neighbor closer to nominal zone than zone $Z$ in the inward direction. The inward is a unit vector directed
from $Z=(X,Y)$ towards the nominal zone $(0,Y)$.
$Z_{\mathrm{ID}}$ the diagonal neighbor in the resultant inward-upstream direction.
$Z_{\mathrm{ON}}$ the neighbor farther from nominal zone than zone $Z$ in the outward direction. The outward is a unit vector
opposite to inward, directed laterally away from the nominal zone $(0,Y)$.
$\alpha_{Z}$ the path consisting $S$ cells that connects node $\widetilde{Z}$ with node $\widetilde{Z}_{\mathrm{UP}}$ in the upstream direction.
$\beta_{Z}$ the path consisting $L$ cells that connects the boundary cell of ${Z}_{\mathrm{IN}}$ with node $\widetilde{Z}$ in the outward direction.
$\gamma_{Z}$ the path consisting $L$ cells that connects node $\widetilde{Z}$ with the boundary cell of ${Z}_{\mathrm{ON}}$ in the outward direction.
$\widehat{\beta}_{Z}$ the set of $L$ diagonal parallel paths that one-one connect cells in $\beta_{Z}$ path with cells of $\alpha_{Z}$ path in the
outward-upstream direction.
$\widehat{\gamma}_{Z}$ the set of $L$ diagonal parallel paths that one-one connect cells in $\gamma_{Z}$ path with cells of $\alpha_{Z}$ path in the
inward-upstream direction.
$\widehat{\delta}_{Z}$ the diagonal path consisting $S$ cells that connects the node $\widetilde{Z}$ with node $\widetilde{Z}_{\mathrm{ID}}$ in the inward-upstream direction.
The UTM provides a heading-reference directed graph to the UAS present in an arbitrary grid level, as shown in Fig. 5(a). For the zone $Z=(X,Y),\ X\neq 0$, the graph comprises $\alpha_{Z},\beta_{Z},\gamma_{Z},\widehat{\beta}_{Z},\widehat{\gamma}_{Z}$ and $\widehat{\delta}_{Z}$ paths (notation described in Table. 1). The UAS motion is constrained by this heading-reference graph. For minimal time delays in the flight path, it is preferable that the UAS transition upstream as close as possible to the nominal segment. When UAS is at node $\widetilde{Z}$ and $Z_{\mathrm{ID}}$ is uncongested, the $\widehat{\delta}_{Z}$ path is available for the UAS to shift segments and move closer to the nominal segment. When $Z_{\mathrm{UP}}$ is uncongested, the $\alpha_{Z}$ path is available to transition upstream. In case both $Z_{\mathrm{UP}}$ and $Z_{\mathrm{ID}}$ are congested, the $\beta_{Z}\rightarrow\gamma_{Z}$ path ($\beta_{Z}$ path $\rightarrow$ node $\widetilde{Z}\rightarrow\gamma_{Z}$ path, collectively) is the alternative path available for UAS to transition outward and avoid the congested zones. If the $Z_{\mathrm{UP}}$ became uncongested while the UAS was in $\beta_{Z}$ or $\gamma_{Z}$ path, then $\widehat{\beta}_{Z}$ or $\widehat{\gamma}_{Z}$ paths are respectively available for UAS to revert towards $Z_{\mathrm{UP}}$.
Figure 5: a) Heading reference directed graph for UAS present in zone $Z$ in an arbitrary level $Y$. b) The $L=5$ slot look-ahead predicted positions of UAS, which are restricted on the heading reference graph. If $M=2$, then by Definition 2, in the perspective of UAS present at node $\widetilde{Z}$, both zones $Z_{\mathrm{UP}}$ and $Z_{\mathrm{ID}}$ are congested (due to UAS 3, 4, 5, and 7). UAS present on $\beta\rightarrow\gamma$ path (UAS 1 and 2) do not contribute to congestion in $Z_{\mathrm{UP}}$ and $Z_{\mathrm{ID}}$.
As a special case, the graph for Stream$(0)$ zones comprises an $\alpha_{0}$ upstream path, two $\gamma_{0}$ paths with opposite headings directed laterally outward from $(0,Y)$ node, and $2L$ number of $\widehat{\gamma}_{0}$ inward diagonal paths. For notation convenience, we use $\alpha_{.},\beta_{.},\gamma_{.},\widehat{\beta}_{.},\widehat{\gamma}_{.},\widehat{\delta}_{.}$, with subscript .$=\mathrm{I},\mathrm{P},\mathrm{O},\mathrm{D}$ referring to the paths in the zones $Z_{\mathrm{IN}},Z_{\mathrm{UP}},Z_{\mathrm{ON}},Z_{\mathrm{ID}}$, respectively. We ignore the subscript when referring to an arbitrary zone.
The rule-based congestion mitigation strategy for the UAS foreseeing congestion in its $Z_{\mathrm{UP}}$ and $Z_{\mathrm{ID}}$ zones is as follows. The following rules assume the UAS is transitioning in a Stream$(X)$ zone, $X=\{0,...,X_{e}\}$. The heading reference graph is symmetric about the nominal segment; hence similar rules when $X=\{-X_{e},...,-1\}$ do hold.
Rule 1.
The UAS present in zone $Z=(X,Y)$ identify respective $Z_{\mathrm{UP}}$ and $Z_{\mathrm{ID}}$ zones as congested or uncongested following Definition 2.
The UAS is restricted to transition on the heading-constrained graph. Under this structural constraint, among the UAS present in the zone neighborhood, the $L$ slot look-ahead predicted positions of only those UAS transitioning on $\alpha_{Z},\widehat{\beta}_{Z},\widehat{\gamma}_{Z},\widehat{\delta}_{\mathrm{O}}$ paths lie in zone $Z_{\mathrm{UP}}$ and are responsible for congestion in $Z_{\mathrm{UP}}$. Similarly, predicted positions of UAS transitioning on $\widehat{\delta}_{Z},\widehat{\beta}_{\mathrm{I}},\widehat{\gamma}_{\mathrm{I}}$ path lie in zone $Z_{\mathrm{ID}}$ (refer Fig. 5(b)).
Rule 2.
In a given timeslot, if the UAS present in $\beta_{Z}\rightarrow\gamma_{Z}$ path finds zone $Z_{\mathrm{UP}}$ congested, then the UAS executes one lateral transition in the outward direction.
In Fig. 6(a), if $Z_{\mathrm{UP}}$ is congested in timeslot $k$, then in timeslot $k+1$ UAS in positions $1,6,9$ would be in positions $2,7,10,$ respectively. A zone congested at one timeslot may or may not be congested in the next. Only myopic position predictions ($1$ slot look-ahead) are possible for UAS transitioning on $\beta\rightarrow\gamma$, as it is uncertain whether these UAS would continue transitioning in $\beta\rightarrow\gamma$ direction or not. Hence, the predicted positions of UAS in ${\beta_{.}}\rightarrow{\gamma_{.}}$, where .$=Z,\mathrm{I},\mathrm{P},\mathrm{O},\mathrm{D}$, are excluded in Definition 2. If a UAS that has entered the $\beta_{Z}$ path (UAS in position $1$ in Fig. 6(a)) finds the $Z_{\mathrm{UP}}$ zone congested for $L$ successive slots, it reaches node $\widetilde{Z}$. Similarly, if the UAS that has entered the $\gamma_{Z}$ path (UAS in position $7$ in Fig. 6(a)) finds the $Z_{\mathrm{UP}}$ zone congested for $L$ successive slots, the UAS ends in the $\beta_{\mathrm{O}}$ path.
Figure 6: a) The relative position $j=\{1,...,S\}$ in the $\beta_{Z}\rightarrow\gamma_{Z}$ path is a cell index with respect to a fixed reference frame in the zone $Z$, where the first cell in $\beta_{Z}$ path, the node $\widetilde{Z}$, and the last cell in $\gamma_{Z}$ path are indexed $j=1,L+1,S$, respectively. The UAS preference $\psi$ to move towards $Z_{\mathrm{UP}}$ varies with its relative position $j$. (Here, $L=5$ slots, $S=2L+1=11$ timeslots). b) $S$ paths connecting cells in the $\beta_{Z}\rightarrow\gamma_{Z}$ path with $\widetilde{Z}_{\mathrm{UP}}$ node.
If $Z_{\mathrm{UP}}$ becomes uncongested in timeslot $k$, then any two UAS in the $\beta_{Z}\rightarrow\gamma_{Z}$ path that simultaneously move towards $Z_{\mathrm{UP}}$ would experience conflict. For example, in Fig. 6(a), UAS in position $2,6,10$ simultaneously reach position ‘d’ by timeslot $k+4$ and conflict with each other. The above risk is overcome by assigning priorities to UAS present in the $\beta_{Z}\rightarrow\gamma_{Z}$ path. It is preferable to have self-enforcing priorities from the UAS perspective. The average number of UAS present in a congested zone and the duration for which the zone remains congested are both measures of congestion. A UAS in the downstream of node $\widetilde{Z}$ has no prior information about congestion in $Z_{\mathrm{UP}}$ as it is unaffected by congestion until it reaches the node (position $6$ in Fig. 6(a)). However, on reaching $\widetilde{Z}$, each timeslot the UAS sees congestion in $Z_{\mathrm{UP}}$, the UAS executes an outward transition on $\gamma_{Z}$ path, and accrues information about congestion in $Z_{\mathrm{UP}}$. Based on this information, the UAS expects the congestion duration in $Z_{\mathrm{UP}}$ to be at least greater than the time interval it has executed outward transitions. Hence, the UAS that has transitioned longer on the $\gamma_{Z}$ path rather prefers to avoid $Z_{\mathrm{UP}}$ and progress towards $Z_{\mathrm{ON}}$ as it expects longer congestion durations in $Z_{\mathrm{UP}}$. In contrast, the UAS that enter $\beta_{Z}$ path (UAS in position $1$ in Fig. 6(a)) have previously avoided $Z_{\mathrm{ID}}$ and have incurred delays, hoping to find $Z_{\mathrm{UP}}$ uncongested. Thus, if $Z_{\mathrm{UP}}$ is uncongested, the UAS in position $1$ actively prefers to move towards $Z_{\mathrm{UP}}$ in comparison to a UAS in the $\gamma_{Z}$ path. Within the zone $Z$, as UAS transitions on $\beta_{Z}\rightarrow\gamma_{Z}$ path, the UAS preference (denoted with $\psi_{Z}$) to move towards $Z_{\mathrm{UP}}$ linearly decreases with each successive outward transition and is modeled as
$\psi_{Z}=L+1-j,\ \ \ j\in\mathbb{Z}_{[1,S]}$ (2)
where $j$ is the relative position of UAS in the $\beta_{Z}\rightarrow\gamma_{Z}$ path. The variation of $\psi$ with relative position $j$ is shown in Fig. 6(a). If a UAS has zero $\psi_{Z}$, it would mean the UAS is currently at node $\widetilde{Z}$. The UAS can execute upstream transitions to reach $\widetilde{Z}_{\mathrm{UP}}$. The UAS with positive (negative) $\psi_{Z}$ can execute outward (inward) - upstream transitions to reach the $Z_{\mathrm{UP}}$ zone. The zone geometry and positional information of the UAS present in the zone are known to every other UAS in the zone. The UAS can compute its preference as well as that of other UAS present in $\beta_{Z}\rightarrow\gamma_{Z}$ using Eqn. 2, enabling the UAS to distributedly decide which UAS should proceed upstream when the respective $Z_{\mathrm{UP}}$ zone becomes uncongested.
Rule 3.
(Priority) If in a given timeslot the $Z_{\mathrm{UP}}$ zone becomes uncongested, then among the UAS transitioning on $\beta_{Z}\rightarrow\gamma_{Z}$ path, that UAS which has the highest preference $\psi_{max}\in\mathbb{Z}_{[-L,L]}$ would proceed towards the $Z_{\mathrm{UP}}$ zone. The remaining UAS execute one outward transition on respective paths.
Applying Rule 3 in Fig. 6(a), if UAS are present in position $1,6,9$, then UAS in position $1$ proceeds to position ‘a’, whereas UAS in position $6,9$ transition to position $7,10,$ respectively. The UAS in position $6,7$ proceed to position ‘b’ if no UAS are present in positions $1$-$5$ and $1$-$6$, respectively. For each relative position $j\in\mathbb{Z}_{[1,S]}$, we have a piecewise re-route path (a sequence of the number of transitions in a specific heading direction) that one-one connects the cells in $\beta_{Z}\rightarrow\gamma_{Z}$ path with node $\widetilde{Z}_{\mathrm{UP}}$. Thus, a total of $2L+1=S$ re-route paths exists. Let $J\subseteq\mathbb{Z}_{[1,S]}$ be the set of positions in the $\beta_{Z}\rightarrow\gamma_{Z}$ path where UAS are present.
Rule 4.
Following Rule 3, the UAS in the $\beta_{Z}\rightarrow\gamma_{Z}$ path that has the highest preference opts the below piecewise re-route path to reach $\widetilde{Z}_{\mathrm{UP}}$ node.
$\psi_{\max}=\max\{\psi_{Z}(j)\ |\ j\in J\}$ (3)
@@
    If $\psi_{\max}=0$, it means the UAS present at node $\widetilde{Z}$ has the highest preference. The UAS executes $S$ successive transitions in upstream direction and reaches $\widetilde{Z}_{\mathrm{UP}}$.
@@
    When $\psi_{max}>0$, it means the UAS in the $\beta_{Z}$ path has the highest preference. The UAS executes $|\psi_{\max}|$ successive transitions in outward-upstream direction followed by $S-|\psi_{\max}|$ upstream transitions and reaches $\widetilde{Z}_{\mathrm{UP}}$.
@@
    When $\psi_{\max}<0$, it means the UAS in the $\gamma_{Z}$ path has the highest preference. The UAS executes $|\psi_{\max}|$ successive transitions in the inward-upstream direction followed by $S-|\psi_{\max}|$ upstream transitions and reaches $\widetilde{Z}_{\mathrm{UP}}$.
Notice that irrespective of the UAS’s position $j$ in $\beta_{Z}\rightarrow\gamma_{Z}$ path, the UAS takes $|\psi_{Z}(j)|+(S-|\psi_{Z}(j)|)=S$ transitions, that is $S$ timeslots to reach the $\widetilde{Z}_{\mathrm{UP}}$ node (refer Fig. 6(b)).
The following rules are for handling contingencies and exceptional scenarios.
Rule 5.
(Descend condition) Consider a UAS transitioning upstream towards node $\widetilde{Z}$. On reaching node $\widetilde{Z}$, if the UAS finds the $Z_{\mathrm{ID}}$ zone uncongested, with no other UAS present in $\beta_{\mathrm{I}}\rightarrow\gamma_{\mathrm{I}}$ path that could possibly move towards $Z_{\mathrm{ID}}$, then the UAS at node $\widetilde{Z}$ executes $S$ successive transitions on $\widehat{\delta}_{Z}$ path and reaches the $\widetilde{Z}_{\mathrm{ID}}$ node.
Here, by descend, we mean the UAS shifts from the higher stream segment (the present parallel segment containing node $\widetilde{Z}$) to a lower stream segment (segment closer to the nominal segment containing node $\widetilde{Z}_{\mathrm{ID}}$). While most conflict risks have been mitigated by imposing priority and descend condition, there still is a rare possibility that two UAS transitioning in the outward and upstream direction simultaneously reach the $\widetilde{Z}$ node in some timeslot $k$ and encounter conflict. Further, following rules 2-4, these UAS will thereafter always conflict with each other for the same cell occupancy. The conflict is resolved as follows.
Rule 6.
(Conflict resolution) If a UAS in the $\beta_{Z}$ path is in conflict with a UAS heading upstream at node $\widetilde{Z}$ (refer Fig. 7), then irrespective of congestion in $Z_{\mathrm{UP}}$, the UAS in the $\beta_{Z}$ path is directed onto the $\alpha_{Z}$ path. The UAS executes $S-1$ transitions in upstream direction and reaches $\widetilde{Z}_{\mathrm{UP}}$.
@@
    if descend condition is satisfied: the UAS heading upstream is directed onto $\widehat{\delta}_{Z}$ path. It executes $S-1$ transitions in inward-upstream direction and reaches $\widetilde{Z}_{\mathrm{ID}}$.
@@
    if descend condition not satisfied: The UAS heading upstream is directed onto the $\gamma_{Z}$ path and executes one lateral transition in outward direction.
@@
The UAS is enclosed in a virtual cylindrical disk of a predefined safety radius. The cells are large enough that during the conflict resolution maneuver, the disks do not overlap. However, the emphasis is on the UAS encountering conflict in time instant $(k-1)\Delta T$ should reach the destined cell centers by time instant $(k+1)\Delta T$. The UAS is free to accelerate or decelerate about the nominal velocity to achieve the above time synchronization. In both Rule 5 and 6, at timeslot $k$ when the descend condition is satisfied or there is a conflict being resolved at node $\widetilde{Z}$, the remaining UAS in the $\beta_{Z}\rightarrow\gamma_{Z}$ path execute one transition in the outward direction, to avoid any conflict risk with the UAS proceeding towards $\widetilde{Z}_{\mathrm{UP}}$ and $\widetilde{Z}_{\mathrm{ID}}$.